OpenStudy (anonymous):

Show that the curve traced parametrically by \(x(t) = (cos(t - 1), t^3 - 1,\frac{1}{t} - 2)\) is tangent to the surface \(x^3 + y^3 + z^3 - xyz = 0 \) when t = 1.

hartnn (hartnn):

could u find \(\Delta z\) where z = x^2+y^3..... = 0 ?

hartnn (hartnn):

the idea is that x(t) will be tangent at t=2 when \(\huge x'(2).\Delta z|_{t=2}=0\)

OpenStudy (anonymous):

May I know what that delta z is?

OpenStudy (anonymous):

And where does that "z = x^2+y^3..... = 0" come from?

hartnn (hartnn):

i should have used other variable, \(\Large w=x^3 + y^3 + z^3 - xyz = 0\)

hartnn (hartnn):

\(\Delta \) is the gradient function, heard of it ?

OpenStudy (anonymous):

I'm sorry... I use \(\nabla\) instead of \(\Delta\), that's why I was confused.

hartnn (hartnn):

eh, i don't remember what was used, you may be correct, so could u find gradient ?

OpenStudy (anonymous):

I can find ∇f, but do I have to substitute values into ∇ f?

hartnn (hartnn):

when t=2, x(2) = ...?

OpenStudy (anonymous):

The problem is ∇f is in terms of x,y,z, so I don't know what to substitute there...

OpenStudy (anonymous):

t=2 or t=1?

hartnn (hartnn):

whatever 3 components u get for t=1 in x(t)

hartnn (hartnn):

i guess, 1,0,-1 so put x=1,y=0 and z=-1

hartnn (hartnn):

oh god, i used t=2 all the time

hartnn (hartnn):

\(\huge x'(1).∇ w|_{x=1,y=-,z=-1}=0\)

hartnn (hartnn):

y=0 i meant

hartnn (hartnn):

0,3,-1 and 3,1,3 dot product comes out to be 0 ? if yes, then tangent

OpenStudy (anonymous):

Why would we find the dot product of x'(1) and ∇w(1,0,-1)?

hartnn (hartnn):

the gradient is orthogonal to the surface at each point. So, if x is tangent to the surface at some point, it should be orthogonal to the gradient of the level surface at that point.

OpenStudy (anonymous):

Thanks a lot :)